OpenStudy (anonymous):

1. Solve S = 4v2 for v

OpenStudy (anonymous):

\[s=4v^2\]\[\frac{s}{4}=v^2\]\[\pm \sqrt{\frac{s}{4}}=v\]

OpenStudy (anonymous):

solve S=4v squared for V